OpenStudy (anonymous):

calculate the percent composition by mass of each element in K2CrO4

OpenStudy (aaronq):

\(\sf mass ~\% ~of ~K =\dfrac{2*Molar~ mass ~of ~K }{Molar ~of~K_2CrO_4 }*100\%\) do this for all the elements. you'll have to use the multiply

OpenStudy (aaronq):

you'll have to mulitply the molar mass by the number of atoms of that element in the compound, like i did with K

OpenStudy (anonymous):

so... mass % of K= 2 mol/195.9961g*100 and that equals to 1.02?

OpenStudy (anonymous):

im not sure if i did that correctly

OpenStudy (aaronq):

you forgot the molar mass of K on the numerator

OpenStudy (anonymous):

is the answer 200?

OpenStudy (aaronq):

200 % ?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

no that can't be. it has to be less than 100% okay, what is the molar mass of K, and what is the molar mass of the compound?

OpenStudy (anonymous):

nevermind is it 39.9%?

OpenStudy (aaronq):

nope

OpenStudy (anonymous):

what...

OpenStudy (aaronq):

\(\sf mass~\% ~of~K=\dfrac{ 2*39}{294}*100\%=27\%\)

OpenStudy (anonymous):

but the molar mass is 194.1903

OpenStudy (joannablackwelder):

You're right, @jleeeh :)

OpenStudy (aaronq):

oh man, i was using the molar mass of potassium dichromate. sorry @jleeeh

OpenStudy (anonymous):

K2CrO4 = 194.1903 g/mol K2 = 78.1966 g/mol Cr = 51.9961 g/mol O4 = 63.9976 g/mol (78.1966 g/mol) / (194.1903 g/mol) = 0.402680 = 40.2680% K (51.9961 g/mol) / (194.1903 g/mol) = 0.267758 = 26.7758% Cr (63.9976 g/mol) / (194.1903 g/mol) = 0.329561 = 32.9561% O